An electron that is moving at a large velocity that is close to the speed of light, has a gamma equal to 5.4.  Calculate the kinetic energy of the electron, in units of Joules.
Answer #1

Relativistic kinetic energy is - K = (C-1) moc? here r=5.4 mo=9.LX10 0-31 kg 2 1 K= (S.4-1 )x9:1810-31 X (3x108) 54.4x9.1X9 X